OpenStudy (anonymous):

1/2+x=5/4 which one ...{1/2,3/4,1,5/4}?

jimthompson5910 (jim_thompson5910):

1/2+x=5/4 x=5/4-1/2 x=5/4-2/4 x=(5-2)/4 x = 3/4